9k=10-k,-1 is this a solution of the equation
Answer by rapaljer(4667)

Substitute -1 in place of the k.

9k = 10 - k
9(-1) = 10 - (-1)
-9 = 10+1

Does it check? What do you think? Is it a solution???

R^2 at SCC